05 September 2009 LBC Checkout

Observers: O. Kuhn, J. Hill (from Tucson)
Telescope Operator: D. Gonzalez Huerta

Summary

Lost to weather.

Preparations

  • TCS Build 32A (everything new, except old PCS kernel)
  • Running IDL on lbtmu06 because of recurring license server problems
  • Aliased "ssh" to "ssh -Y" for user LBTO on lbtdu16 to cure Xdisplay problems
  • Reduced DOFPIA convergence threshold for Z22 from 100 to 50, as we think 100 nm is still a large amount of uncorrected Z22.

Details

Thunderstorms and light rain early stopped us from opening. We threatened to open a few times, but never did.

We tested Stephen's new copydown.sh script, and improved it with file size checking.

-- JohnHill - 05 Sep 2009
Topic revision: r2 - 05 Sep 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback